Are the following lines parallel, perpendicular, or neither?

L1 through (-4,-7) and (1,3)
L2 through (2,6) and (4,10)

Answer by Nate(3500)

M = (y2 - y1)/(x2 - x1) = (3 + 7)/(1 + 4) = 10/5 = 2
M = (y2 - y1)/(x2 - x1) = (10 - 6)/(4 - 2) = 4/2 = 2
The lines are parallel.